User name cusanno
Log entry time 08:03:59 on November 9,2006
Entry number 189135
We keep running on Carbon target as on previous shift.
Bob Michaels is still perfoming tests with Compton.
00:30 MCC called reporting some problems on IWL cooling system at injector. They are worring about our target, but target is OK.
01:16 Right Arm Q1 tripped. Related to the IWL problem?
Bob Michaels take chance to get nearly one hour of Compton test run.
Run 2037 and 2038 are Qasy tests.
02:00 Right Arm Q1 recovered
02:11 Bob Michaels ended Compton tests, we go back to 85uA
02:40 Left Dipole IOC frozen, rebooted. Ended run 2038.
02:55 start new run 2039.
03:30 Left Dipole IOC frozen again, ended run 2040.
Rebooted. Start new run 2041
7:40 Data taking ended, asked for Access.
Production Run on Carbon:
Test run for QASY: